OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 4, 12, 36, 108, ... Converges; 484 Diverges Converges; 52 Converges; 160 @ganeshie8 @phi @beccaboo333

OpenStudy (anonymous):

will give best answer & m edal!

OpenStudy (phi):

if you divide 12/4 you get 3 if you divide 36/12 you get 3 if you divide 108/36 you get 3

OpenStudy (kainui):

First off, have you tried to find the pattern here? It helps to decompose each term into its prime factors and you should be able to notice a pattern.

OpenStudy (anonymous):

yes @phi

OpenStudy (phi):

what do you think the next number is after 108 ?

OpenStudy (anonymous):

324

OpenStudy (phi):

and the number after 324 ?

OpenStudy (anonymous):

972

OpenStudy (phi):

any chance the next number is "closing in" on some number or is it "diverging" ?

OpenStudy (anonymous):

i think its closing in on a number

OpenStudy (anonymous):

or is it diverginng?

OpenStudy (phi):

I guess you don't know the definitions these words. The next number is 3 * the previous number. The next number is bigger than the previous if we keep going, we get bigger and bigger numbers. If we go far enough we get gigantic enormous humongous numbers.

OpenStudy (anonymous):

yes, does that mean it diverges because it doesn't stop at a certain number? or is it converge because its getting bigger?

OpenStudy (phi):

diverge means we end up with gigantic enormous humongous numbers.

OpenStudy (anonymous):

Okay, thank u!

OpenStudy (phi):

converge means the "next number" is getting closer to some number example: 1/x when x =1 we get 1/1 =1 when x = 10 we get 1/10 = 0.1 x=100 , we get 1/100 = 0.01 x=1000000 , we get 0.0000001 get are getting closer and closer to 0 we are converging.

OpenStudy (anonymous):

ohhh okay, understood. thank u so much :)

OpenStudy (anonymous):

\[\text{Table}\left[4\text{ * }3^{n-1},\{n,1,10\}\right]= \]{4, 12, 36, 108, 324, 972, 2916, 8748, 26244, 78732}